# Kazue Kudo

### v4

@@ I Introduction
This paper presents a method for solving PDEs using an Ising machine. Ising machines are special-purpose computers designed to efficiently solve combinatorial optimization problems. They have been realized using various technologies, including a quantum annealer [12], digital processors based on simulated annealing (SA) [13, 14, 15, 16, 17, 18, 19], those based on simulated bifurcation [20, 21, 22, 23], and coherent Ising machines [24, 25, 26, 27, 28]. Although initially designed to solve combinatorial optimization problems, they have applications in various fields. This work demonstrates their potential for solving PDEs and expands their application fields.
For solving a problem using an Ising machine, the problem is represented by the quadratic unconstrained binary optimization (QUBO) formulation. In the QUBO formulation, the solution to a PDE is expressed as a linear combination of binary variables. The number of binary variables, which depends on the desired precision and the discretization mesh size, is a critical factor in QUBO-based methods. This work employs an iterative method, inspired by Ref. [11], that uses a relatively small number of variables. In a usual QUBO formulation, increasing precision typically requires more binary variables. However, the proposed algorithm enables the computation of eigenvectors to arbitrary precision without increasing the number of variables, although more iterations are required. Another advantage of the algorithm is that the desired accuracy becomes available even if the performance of the annealing procedure is imperfect. This paper introduces a modified algorithm that improves upon the original one and proposes a more efficient precision update scheme.
@@ II-A Generalized eigenvalue problem
We consider the SLE arising from the discretization of a PDE:
@@
where $K$ is a positive-definite matrix, and $\bm{u}$ and $\bm{f}$ are real vectors. While SLEs can be defined in complex spaces, we focus on real-valued systems for simplicity. The goal is to determine the solution $\bm{u}$ for given $K$ and $\bm{f}$.
@@ II-B Algorithm
where $A,B\in\mathbb{R}^{n\times n}$ are symmetric matrices, and $B\in\mathbb{R}^{n\times n}$ is positive definite. The eigenvector $\bm{v}\in\mathbb{R}^{n}$ is approximated as $\bm{v}=(I_{n}\otimes\bm{p})\bm{q}$, where $\bm{p}=\left(-1,\frac{1}{2},\frac{1}{2^{2}},\ldots,\frac{1}{2^{b-1}}\right)$ is a row vector, and $\bm{q}=\{0,1\}^{nb}$ is a binary vector. The parameter $b$ is determines the precision of the initial guess. The explicit form of $\bm{v}$ is written as
@@ II-B2 Iterative descent stage
The procedure in the iterative descent stage is summarized as follows. First, set the initial mesh size: $r=r_{\rm init}$. Then, the following procedure is repeated until $r$ reaches to the tolerance precision $\epsilon_{0}$ or the eivenvalue $\lambda$ converges.
@@ III-A Dependence on precision
At the initial guess stage, as shown in Fig. 1(a), the number of iterations is relatively insensitive to the precision parameter $b$. The number of iterations is smaller for a shorter annealing time, and it decreases as the system size increases. The standard deviation, which is not depicted, typically ranges from 0.5 to 1.0.
@@ III-B Dependence on system size
Figure 4: The average number of iterations at the iterative descent stage for solving the one-dimensional symmetric, asymmetric, and two-dimensional Poisson equations as a function of system size $n$. Results are shown for $b=5$ and $N_{\rm step}=10^{3},\,10^{4}$. Symbol size is proportional to the success rate. The dashed line represents $ce^{\alpha n}$ with constants $c$ and $\alpha$.
As shown in the above, the number of iterations depends on system size and annealing time. Figure 4 shows the average number of iterations required at the iterative descent stage for solving the one-dimensional symmetric, asymmetric, and two-dimensional Poisson equations as a function of system size $n$. The precision parameter is fixed as $b=5$, and results for $N_{\rm step}=10^{3}$ and $10^{4}$ are shown, where the lower one of two points with the same symbol type at the same $n$ corresponds to $N_{\rm step}=10^{4}$. Note that the symbol size is proportional to the success rate, and several points are too small to recognize, particularly for larger system sizes. The dashed line, representing an exponential function $ce^{\alpha n}$ with $c=13$ and $\alpha=0.03$, provides a reference for the growth rate. For the one-dimensional symmetric and two-dimensional Poisson equations, the increase in iterations is slower than exponential.
@@ IV Discussion
The error evaluated to calculate the success rate was based on the numerical solution of the SLE. However, the underlying goal is to solve the original PDE. The discretization error, which is inherent to numerical methods, plays a crucial role in determining the overall accuracy of the solution. In the case of central difference discretization, the error is typically of order $O(h^{2})$, where $h=n+1$ is the mesh size. To improve the accuracy of the PDE solution, it is necessary to increase $n$, which in turn increases the computational cost.
@@
Traditional numerical methods for solving eigenvalue problems typically have a computational complexity of $O(n^{3})$. In contrast, the proposed method exhibits exponential or subexponential growth in computational time with increasing system size, as illustrated in Fig. 4. However, the coefficient $\alpha$ in the exponential function $e^{\alpha n}$ is relatively small, even for the asymmetric Poisson equation. This suggests the proposed method, together with using an Ising machine, can outperform conventional methods for certain classes of simple problems. It is important to note that this analysis focuses on the number of iterations rather than the actual annealing time. In the simulated annealing method used in this work, the number of sampling steps is proportional to $nb$. Moreover, solving large-scale problems often requires longer annealing times, as discussed earlier.

### v5

@@ I Introduction
This paper presents a method for solving PDEs using an Ising machine. Ising machines are special-purpose computers designed to efficiently solve combinatorial optimization problems. They have been realized using various technologies, including a quantum annealer [12], digital processors based on simulated annealing (SA) [13, 14, 15, 16, 17], those based on simulated bifurcation [18, 19, 20, 21], and coherent Ising machines [22, 23, 24, 25, 26]. Although initially designed to solve combinatorial optimization problems, they have applications in various fields. This work demonstrates their potential for solving PDEs and expands their application fields.
For solving a problem using an Ising machine, the problem is represented by the quadratic unconstrained binary optimization (QUBO) formulation. In the QUBO formulation, the solution to a PDE is expressed as a linear combination of binary variables. The number of binary variables, which depends on the desired precision and the discretization mesh size, is a critical factor in QUBO-based methods. This work employs an iterative method, inspired by Ref. [11], that uses a relatively small number of variables. In a usual QUBO formulation, increasing precision typically requires more binary variables. However, the proposed algorithm enables the computation of eigenvectors to arbitrary precision without increasing the number of variables. The key idea is to improve accuracy by refining the mesh-size scale while keeping the number of divisions fixed. Another advantage of the algorithm is that the desired accuracy becomes available even if the performance of the annealing procedure is imperfect. This paper introduces a modified algorithm that improves upon the original one and proposes a more efficient precision update scheme.
@@ II-A Generalized eigenvalue problem
We consider the SLE arising from the discretization of a PDE with Dirichlet boundary conditions:
@@
where $K$ is a symmetric positive-definite matrix, and $\bm{u}$ and $\bm{f}$ are real vectors. While SLEs can be defined in complex spaces, we focus on real-valued systems for simplicity. The goal is to determine the solution $\bm{u}$ for given $K$ and $\bm{f}$.
@@ II-B Algorithm
where $A,B\in\mathbb{R}^{n\times n}$ are symmetric matrices, and $B\in\mathbb{R}^{n\times n}$ is positive definite. The eigenvector $\bm{v}\in\mathbb{R}^{n}$ is approximated as $\bm{v}=(I_{n}\otimes\bm{p})\bm{q}$, where $\bm{p}=\left(-1,\frac{1}{2},\frac{1}{2^{2}},\ldots,\frac{1}{2^{b-1}}\right)$ is a row vector, and $\bm{q}=\{0,1\}^{nb}$ is a binary vector. The parameter $b$ determines the precision of the initial guess. The explicit form of $\bm{v}$ is written as
@@ II-B2 Iterative descent stage
The procedure in the iterative descent stage is summarized as follows. First, set the initial mesh size: $r=r_{\rm ini}$. Then, the following procedure is repeated until $r$ reaches to the tolerance precision $\epsilon_{0}$ or the eigenvalue $\lambda$ converges.
@@ III Results
In this study, we performed SA using the open-source Python library OpenJij [28]. The SA sampler in this library automatically determines a geometric temperature schedule.
@@ III-A Dependence on precision
At the initial guess stage, as shown in Fig. 1(a), the number of iterations is relatively insensitive to the precision parameter $b$. The standard deviation, which is not depicted, typically ranges from 0.5 to 1.0. The number of iterations is smaller for shorter annealing times and decreases as the system size increases. These trends are different from those in Figs. 1(b) or (c), which reflects the difference in objective functions.
@@ III-B Dependence on system size
Figure 4: The average number of iterations at the iterative descent stage for solving the one-dimensional symmetric, one-dimensional symmetric asymmetric, and two-dimensional Poisson equations as a function of system size $n$. Results for $N_{\rm step}=10^{3},\,10^{4}$ are shown for (a) $b=3$ and (b) $b=5$. Symbol size is proportional to the success rate. The dashed line represents $ce^{\alpha n}$ with constants $c$ and $\alpha$.
As shown in the above, the number of iterations depends on system size and annealing time. Figure 4 shows the average number of iterations required at the iterative descent stage for solving the one-dimensional symmetric, one-dimensional asymmetric, and two-dimensional Poisson equations as a function of system size $n$. The precision parameters are $b=3$ and $b=5$ in Figs. 4(a) and 4(b), respectively. Results for $N_{\rm step}=10^{3}$ and $10^{4}$ are shown in each panel, where the lower one of two points with the same symbol type at the same $n$ corresponds to $N_{\rm step}=10^{4}$. Note that the symbol size is proportional to the success rate, and several points are too small to recognize, particularly for larger system sizes. The dashed line, representing an exponential function $ce^{\alpha n}$ with $c=12$ and $\alpha=0.04$, provides a reference for the growth rate. For the one-dimensional symmetric and two-dimensional Poisson equations, the increase in iterations is slower than exponential.
@@ IV Discussion
The error evaluated to calculate the success rate was based on the numerical solution of the SLE. However, the underlying goal is to solve the original PDE. The discretization error, which is inherent to numerical methods, plays a crucial role in determining the overall accuracy of the solution. In the case of central difference discretization, the error is typically of order $O(h^{2})$, where $h=1/(n+1)$ is the mesh size. To improve the accuracy of the PDE solution, it is necessary to increase $n$, which in turn increases the computational cost.
@@
Linear-solver algorithms for PDEs typically have computational complexities ranging from $O(n)$ to $O(n^{2})$. In contrast, the proposed method exhibits exponential or subexponential growth in computational time with increasing system size, as illustrated in Fig. 4. However, the coefficient $\alpha$ in the exponential function $e^{\alpha n}$ is relatively small, even for the asymmetric Poisson equation. This suggests the proposed method, together with using an Ising machine, can outperform conventional methods for certain classes of simple problems. It is important to note that this analysis focuses on the number of iterations rather than the actual annealing time. In the simulated annealing method used in this work, the number of sampling steps is proportional to $nb$. Moreover, solving large-scale problems often requires longer annealing times, as discussed earlier.